# MoDGS: Dynamic Gaussian Splatting from Casually-captured Monocular Videos

### v2

@@ Abstract
In this paper, we propose MoDGS, a new pipeline to render novel views of dynamic scenes from a casually captured monocular video. Previous monocular dynamic NeRF or Gaussian Splatting methods strongly rely on the rapid movement of input cameras to construct multiview consistency but struggle to reconstruct dynamic scenes on casually captured input videos whose cameras are either static or move slowly. To address this challenging task, MoDGS adopts recent single-view depth estimation methods to guide the learning of the dynamic scene. Then, a novel 3D-aware initialization method is proposed to learn a reasonable deformation field and a new robust depth loss is proposed to guide the learning of dynamic scene geometry. Comprehensive experiments demonstrate that MoDGS is able to render high-quality novel view images of dynamic scenes from just a casually captured monocular video, which outperforms state-of-the-art methods by a significant margin. The code will be publicly available. Project page:   https://MoDGS.github.io
@@ 1 Introduction
Dynamic View Synthesis (DVS) has achieved impressive improvements along with the emerging neural representations (Mildenhall et al., 2020) and Gaussian splatting (Kerbl et al., 2023) techniques. Most of the existing DVS methods (Cao & Johnson, 2023; Yang et al., 2023b) require multiview videos captured by dense synchronized cameras to achieve good rendering quality. Though some works can process a monocular video for DVS, as pointed out by DyCheck (Gao et al., 2022), these methods require the camera of the monocular video to have extremely large movements, which is called “Teleporting Camera Motion” on different viewpoints, so these methods can utilize the multiview consistency provided by this pseudo multiview video to reconstruct the 3D geometry of the dynamic scene. However, such large camera movements are rarely seen in casually captured videos because casual videos are usually produced by smoothly moving or even static cameras. When the camera moves slowly or is static, the multiview consistency constraint will be much weaker and all these existing DVS methods fail to produce high-quality novel-view images, as shown in Fig. 1.
In this paper, we present Monocular Dynamic Gaussian Splatting (MoDGS) to render novel-view images from casually captured monocular videos in a dynamic scene. MoDGS addresses the weak multiview constraint problem by adopting a monocular depth estimation method  (Fu et al., 2024), which provides prior depth information on the input video to help the 3D reconstruction.
@@
However, we find that simply applying a single-view depth estimator in DVS to supervise rendered depth maps is not enough for high-quality novel view synthesis. First, the depth supervision only provides information for each frame but does not help to associate 3D points between two frames in time. Thus, we still have difficulty in learning an accurate time-dependent deformation field. Second, the estimated depth values are not consistent among different frames.
@@
To learn a robust deformation field from a monocular video, we propose a 3D-aware initialization scheme for the deformation field. Existing methods (Katsumata et al., 2023) solely rely on supervision from 2D flow estimation, which produces deteriorated results without sufficient multiview consistency. We find that directly initializing the deformation field in the 3D space greatly helps the subsequent learning of the 4D representations and improves the rendering quality as shown in Fig. 1.
To better utilize the estimated depth maps for supervision, we propose a novel depth loss to address the scale inconsistency of estimated depth values across different frames.
@@
Previous methods (Li et al., 2023b; Liu et al., 2023a) supervise the rendered depth maps using a scale-invariant depth loss by minimizing the $L2$ distance of normalized rendered depth and depth priors, and the most recent method  (Zhu et al., 2023) proposed to supervise the rendered depth maps using a Pearson correlation loss to mitigate the scale ambiguity between the reconstructed scene and the estimated depth maps. However, the estimated depth maps of different frames are not even consistent after normalizing to the same scale. To address these challenges, we observe that despite the inconsistency in values, the orders of depth values of different pixels in different frames are stable, which motivates us to propose an ordinal depth loss. This novel ordinal depth loss enables us to fully utilize the estimated depth maps for high-quality novel view synthesis.
@@
Figure 1: Given a casually captured monocular video of a dynamic scene, MoDGS is able to synthesize high-quality novel-view images in this scene. In the middle column, the baseline method (Yang et al., 2023b) fails to correctly reconstruct the 3D dynamic scenes on this static monocular video. The white regions in cyan bounding boxes are not visible in the input video (red bounding boxes) so there are some artifacts for these invisible regions. In the rightmost column, the input estimated monocular depth is inconsistent (red bounding boxes); however, our proposed ordinal depth loss effectively ensures more consistent depth outputs. This loss enhances the accuracy and reliability of learning underlying geometry.
@@ 2 Related Work
Dynamic Gaussian Splatting. The recent emergence of 3D Gaussian Splatting (3DGS) demonstrates its efficacy for super-fast real-time rendering attributed to its explicit point cloud representation. Recent follow-ups extend 3DGS to model dynamic 3D scenes. Luiten et al. (2023) track dynamic 3D Gaussians by frame-by-frame training from synchronized multi-view videos. Yang et al. (2023b) propose a deformable version of 3DGS by introducing a deformation MLP network to model the 3D flows. Wu et al. (2023) and Duisterhof et al. (2023) also introduce a deformation field but using a more efficient Hexplane representation (Cao & Johnson, 2023). Yang et al. (2023a) proposes a dynamic representation with a collection of 4D Gaussian primitives, where the time evolution can be encoded by 4D spherical harmonics. Bae et al. (2024) encodes motions with a per-Gaussian feature vector. Some other works (Li et al., 2023a; Lin et al., 2023; Liang et al., 2023) also study how to effectively encode the motions for Gaussians with different bases. To effectively learn the motions of Gaussians, some works (Feng et al., 2023; Yu et al., 2023; Huang et al., 2024) resort to clustering the motions together for a compact representation.
@@
The ordinal relation between pixels has been investigated in recent years, especially in the field of monocular depth estimation.  Zoran et al. (2015) proposed to use a three-category classification network to predict the order relation of given pixel pairs. Then the depth can be extracted by optimizing a constrained quadratic optimization problem. Similarly,  Fu et al. (2018) treats the depth prediction problem as a multi-class classification problem.
Chen et al. (2016) further proposes a ranking loss to learn metric depth, which encourages a small difference between depths if the ground-truth relation is equal; otherwise it encourages a large difference. Then, Pavlakos et al. (2018) extends this differentiable ranking loss to the human pose estimation task. However, these works only utilize limited numbers of depth orders for training (one pair in  Chen et al. (2016) and 17 pairs in  Pavlakos et al. (2018)), resulting in coarse supervision for depth maps. The direct application of their ranking loss as depth supervision has yet to be explored. Moreover, our ordinal depth takes the rendered metric depth maps as input, which are dense grids of float numbers. Our task is different from previous depth estimation and pose estimation tasks and we present a comparison between our ordinal depth loss and depth ranking loss in Appendix A.10.
@@ 3.1 Gaussians and Deformation Fields
After getting the initialized deformation field, we will initialize a set of 3D Gaussians in the canonical space as shown in Fig. 3 (right). We achieve this by first converting all the depth maps to get 3D points. Then, these 3D points are deformed backward to the canonical 3D space. This means that we transform all the depth points of all timestamps to the canonical space, which results in a large amount of points. We then evenly downsample these points with a predefined voxel size to reduce the point number and we initialize all our Gaussians with the locations of these downsampled 3D points in the canonical space.
@@ 3.3 Ordinal Depth Loss
Existing dynamic Gaussian Splatting or NeRF methods also adopt a depth loss to supervise the learning of their 3D representations. One possible solution (Zhu et al., 2023; Li et al., 2021; Liu et al., 2023a) is to maximize a Pearson correlation between the rendered depth and the estimated single-view depth
@@ 4 Experiments
Figure 5: Qualitative comparison on the novel-view renderings of the DyNeRF (Li et al., 2022) and Nvidia (Yoon et al., 2020) datasets. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023b), and HexPlane (Cao & Johnson, 2023).
Figure 6: Qualitative comparison of DVS quality on the MCV dataset. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023b), HexPlane (Cao & Johnson, 2023), and RoDynRF (Liu et al., 2023a).
@@ 4.1 Evaluation Protocols
Previous DVS methods (Cao & Johnson, 2023; Yang et al., 2023b; Gao et al., 2021) all use different cameras to train the dynamic NeRF or Gaussian Splatting. Even though they only use one camera at a specific timestamp, they use different cameras at different timestamps so that a pseudo multiview video can be constructed to learn the 3D structures of the scene. Since our target is to conduct novel-view synthesis on the casually captured images, we do not adopt such ”teleporting camera motions” to construct training videos but just adopt one static camera to record training videos. Then, we render the images from the viewpoints of another camera for evaluation.
@@
We compare MoDGS with 4 baseline methods to demonstrate the superior ability of MoDGS to synthesize novel-view images with casually captured monocular videos. These methods can be categorized into two classes. The first is the NeRF-based methods including HexPlane (Cao & Johnson, 2023) and RoDynRF (Liu et al., 2023a). HexPlane represents the scene with six feature planes in both 3D space and time-space. We find that HexPlane does not produce reasonable results if only a monocular video from a single camera is given as input. Thus, other than the input monocular video, we use another video from a different viewpoint to train HexPlane for the DVS task. RoDynRF is a SoTA NeRF-based DVS method that also adopts single-view depth estimation as supervision for the 3D dynamic representations. We train it with the same single-view depth estimator GeoWizard (Fu et al., 2024) as ours. The second class is the Gaussian Splatting-based DVS methods including Deformable-GS (Yang et al., 2023b) and SC-GS (Huang et al., 2024). Deformable-GS also associates a deformation field with a set of canonical Gaussians for DVS. SC-GS learns a set of keypoints and uses the deformation of these keypoints to blend the deformation of arbitrary 3D points.
@@ 4.4 Limitations
Though our method can conduct dynamic view synthesis from casually captured monocular videos, the task is still extremely challenging. One limitation is that our method can only reconstruct the visible 3D parts but cannot imagine the unseen parts, which leads to artifacts when rendering novel view videos on these unseen parts. Incorporating recent 3D-related diffusion generative models (Chung et al., 2023a; Liu et al., 2023b; Long et al., 2023) could be a promising direction to solve this problem, which we leave for future works. Another limitation is that the current training time is comparable with existing DVS methods, which could take several hours for a single scene. How to efficiently reconstruct the dynamic field would be an interesting and promising future research topic. Meanwhile, when the camera is completely static, our method strongly relies on the single-view depth estimator to estimate the 3D depth maps. Though existing single-view depth estimators (Fu et al., 2024; Ke et al., 2023; Yang et al., 2024; Bhat et al., 2023) are trained on large-scale datasets and predict reasonable depth maps for most cases, these depth estimators may fail to capture some details which degenerate the quality. Meanwhile, specular reflection can not be robustly and well handled, which is also an extremely challenging problem in the monocular setting, which we leave for future works.
@@ 5 Conclusion
In this paper, we have presented a novel dynamic view synthesis paradigm called MoDGS. In comparison with existing DVS methods which requires “teleporting camera motions”, MoDGS is designed to render novel view images from a casually captured monocular video. MoDGS introduces two new designs to finish this challenging task. First, a new 3D-aware initialization scheme is proposed, which directly initializes the deformation field to provide a reasonable starting point for subsequent optimization. We further analyze the problem of depth loss and propose a new ordinal depth loss to supervise the learning of the scene geometry. Extensive experiments on three datasets demonstrate superior performances of our method on in-the-wild monocualr videos over baseline methods.
@@ A.1 Implementation Details
We implement our MoDGS with PyTorch. To initialize the deformation field, we train it with 20k steps as stated in Sec. 3.2. Subsequently, we jointly train the 3D Gaussians and the deformation field with the rendering loss and the ordinal depth loss for another 20k steps. In Sec. 3.2, the flow is computed in evenly sampled key frames($1/5$). And the downsampling voxel size for Gaussian initialization is $0.004^{3}$ (scenes are normalized to $[-1,1]^{3}$ ). For the outer optimization loop and rendering loss, we exactly follow the original 3DGS. And we use Gaussian centers to render depth  (Yang et al., 2023b). We adopt an Adam optimizer for optimization. The learning rate for 3D Gaussians exactly follows the official implementation of 3D GS (Kerbl et al., 2023), while the learning rate of the deformation network undergoes exponential decay from 1e-3 to 1e-4 in initialization and from 1e-4 to 1e-6 in the subsequent optimization. We set $\alpha=100$ for $\ell_{\text{ordinal}}$. The weight of our depth order loss is 0.1. When computing depth ordinal loss, we first normalize the depth range to [0,1] and we only consider the depth pair with a difference larger than 0.02 for loss computation. The whole training takes around 3.5 hours to converge (2 hours for the initialization and 1.5 hours for the subsequent optimization) on an NVIDIA RTX A6000 GPU, which uses about 14G memory. The rendering speed of MoDGS is about 75 FPS.

### v3

@@ Abstract
In this paper, we propose MoDGS, a new pipeline to render novel views of dynamic scenes from a casually captured monocular video. Previous monocular dynamic NeRF or Gaussian Splatting methods strongly rely on the rapid movement of input cameras to construct multiview consistency but struggle to reconstruct dynamic scenes on casually captured input videos whose cameras are either static or move slowly. To address this challenging task, MoDGS adopts recent single-view depth estimation methods to guide the learning of the dynamic scene. Then, a novel 3D-aware initialization method is proposed to learn a reasonable deformation field and a new robust depth loss is proposed to guide the learning of dynamic scene geometry. Comprehensive experiments demonstrate that MoDGS is able to render high-quality novel view images of dynamic scenes from just a casually captured monocular video, which outperforms state-of-the-art methods by a significant margin. Project page:   https://MoDGS.github.io
@@ 1 Introduction
Dynamic View Synthesis (DVS) has achieved impressive improvements along with the emerging neural representations (Mildenhall et al., 2020) and Gaussian splatting (Kerbl et al., 2023) techniques. Most of the existing DVS methods (Cao & Johnson, 2023; Yang et al., 2023) require multiview videos captured by dense synchronized cameras to achieve good rendering quality. Though some works can process a monocular video for DVS, as pointed out by DyCheck (Gao et al., 2022), these methods require the camera of the monocular video to have extremely large movements, which is called “Teleporting Camera Motion” on different viewpoints, so these methods can utilize the multiview consistency provided by this pseudo multiview video to reconstruct the 3D geometry of the dynamic scene. However, such large camera movements are rarely seen in casually captured videos because casual videos are usually produced by smoothly moving or even static cameras. When the camera moves slowly or is static, the multiview consistency constraint will be much weaker and all these existing DVS methods fail to produce high-quality novel-view images, as shown in Fig. 1.
@@
In this paper, we present Monocular Dynamic Gaussian Splatting (MoDGS) to render novel-view images from casually captured monocular videos in a dynamic scene. MoDGS addresses the weak multiview constraint problem by adopting a monocular depth estimation method  (Fu et al., 2024), which provides prior depth information on the input video to help the 3D reconstruction. However, we find that simply applying a single-view depth estimator in DVS to supervise rendered depth maps is not enough for high-quality novel view synthesis. First, the depth supervision only provides information for each frame but does not help to associate 3D points between two frames in time. Thus, we still have difficulty in learning an accurate time-dependent deformation field. Second, the estimated depth values are not consistent among different frames.
@@
To learn a robust deformation field from a monocular video, we propose a 3D-aware initialization scheme for the deformation field. Existing methods (Katsumata et al., 2024) solely rely on supervision from 2D flow estimation, which produces deteriorated results without sufficient multiview consistency. We find that directly initializing the deformation field in the 3D space greatly helps the subsequent learning of the 4D representations and improves the rendering quality as shown in Fig. 1.
@@
To better utilize the estimated depth maps for supervision, we propose a novel depth loss to address the scale inconsistency of estimated depth values across different frames. Previous methods (Li et al., 2023b; Liu et al., 2023a) supervise the rendered depth maps using a scale-invariant depth loss by minimizing the $L2$ distance of normalized rendered depth and depth priors, and the most recent method  (Zhu et al., 2023c) proposed to supervise the rendered depth maps using a Pearson correlation loss to mitigate the scale ambiguity between the reconstructed scene and the estimated depth maps. However, the estimated depth maps of different frames are not even consistent after normalizing to the same scale. To address these challenges, we observe that despite the inconsistency in values, the orders of depth values of different pixels in different frames are stable, which motivates us to propose an ordinal depth loss. This novel ordinal depth loss enables us to fully utilize the estimated depth maps for high-quality novel view synthesis.
@@
Figure 1: Given a casually captured monocular video of a dynamic scene, MoDGS is able to synthesize high-quality novel-view images in this scene. In the middle column, the baseline method (Yang et al., 2023) fails to correctly reconstruct the 3D dynamic scenes on this static monocular video. The white regions in cyan bounding boxes are not visible in the input video (red bounding boxes) so there are some artifacts for these invisible regions. In the rightmost column, the input estimated monocular depth is inconsistent (red bounding boxes); however, our proposed ordinal depth loss effectively ensures more consistent depth outputs. This loss enhances the accuracy and reliability of learning underlying geometry.
@@ 2 Related Work
Dynamic Gaussian Splatting. The recent emergence of 3D Gaussian Splatting (3DGS) demonstrates its efficacy for super-fast real-time rendering attributed to its explicit point cloud representation. Recent follow-ups extend 3DGS to model dynamic 3D scenes. Luiten et al. (2023) track dynamic 3D Gaussians by frame-by-frame training from synchronized multi-view videos. Yang et al. (2023) propose a deformable version of 3DGS by introducing a deformation MLP network to model the 3D flows. Wu et al. (2024) and Duisterhof et al. (2023) also introduce a deformation field but using a more efficient Hexplane representation (Cao & Johnson, 2023). Yang et al. (2024c) proposes a dynamic representation with a collection of 4D Gaussian primitives, where the time evolution can be encoded by 4D spherical harmonics. Bae et al. (2024) encodes motions with a per-Gaussian feature vector. Some other works (Li et al., 2023a; Lin et al., 2024; Liang et al., 2023) also study how to effectively encode the motions for Gaussians with different bases. To effectively learn the motions of Gaussians, some works (Feng et al., 2024; Yu et al., 2023; Huang et al., 2024) resort to clustering the motions together for a compact representation.
@@
The ordinal relation between pixels has been investigated in recent years, especially in the field of monocular depth estimation.  Zoran et al. (2015) proposed to use a three-category classification network to predict the order relation of given pixel pairs. Then the depth can be extracted by optimizing a constrained quadratic optimization problem. Similarly,  Fu et al. (2018) treats the depth prediction problem as a multi-class classification problem. Chen et al. (2016) further proposes a ranking loss to learn metric depth, which encourages a small difference between depths if the ground-truth relation is equal; otherwise it encourages a large difference. Then, Pavlakos et al. (2018) extends this differentiable ranking loss to the human pose estimation task. However, these works only utilize limited numbers of depth orders for training (one pair in  Chen et al. (2016) and 17 pairs in  Pavlakos et al. (2018)), resulting in coarse supervision for depth maps. The direct application of their ranking loss as depth supervision has yet to be explored. Moreover, our ordinal depth takes the rendered metric depth maps as input, which are dense grids of float numbers. Our task is different from previous depth estimation and pose estimation tasks and we present a comparison between our ordinal depth loss and depth ranking loss in Appendix A.10.
@@ 3.1 Gaussians and Deformation Fields
After getting the initialized deformation field, we will initialize a set of 3D Gaussians in the canonical space as shown in Fig. 3 (right). We achieve this by first converting all the depth maps to get 3D points. Then, these 3D points are deformed backward to the canonical 3D space. This means that we transform all the depth points of all timestamps to the canonical space, which results in a large amount of points. We then evenly downsample these points with a predefined voxel size to reduce the point number and we initialize all our Gaussians with the locations of these downsampled 3D points in the canonical space. Here, more advanced learning-based adaptive fusion strategies (You et al., 2023) could be adopted to downsample the points, potentially improving the representation.
@@ 3.3 Ordinal Depth Loss
Existing dynamic Gaussian Splatting or NeRF methods also adopt a depth loss to supervise the learning of their 3D representations. One possible solution (Zhu et al., 2023c; Li et al., 2021; Liu et al., 2023a) is to maximize a Pearson correlation between the rendered depth and the estimated single-view depth
@@ 4 Experiments
Figure 5: Qualitative comparison on the novel-view renderings of the DyNeRF (Li et al., 2022) and Nvidia (Yoon et al., 2020) datasets. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023), and HexPlane (Cao & Johnson, 2023).
Figure 6: Qualitative comparison of DVS quality on the MCV dataset. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023), HexPlane (Cao & Johnson, 2023), and RoDynRF (Liu et al., 2023a).
@@ 4.1 Evaluation Protocols
Previous DVS methods (Cao & Johnson, 2023; Yang et al., 2023; Gao et al., 2021) all use different cameras to train the dynamic NeRF or Gaussian Splatting. Even though they only use one camera at a specific timestamp, they use different cameras at different timestamps so that a pseudo multiview video can be constructed to learn the 3D structures of the scene. Since our target is to conduct novel-view synthesis on the casually captured images, we do not adopt such ”teleporting camera motions” to construct training videos but just adopt one static camera to record training videos. Then, we render the images from the viewpoints of another camera for evaluation.
@@
We compare MoDGS with 4 baseline methods to demonstrate the superior ability of MoDGS to synthesize novel-view images with casually captured monocular videos. These methods can be categorized into two classes. The first is the NeRF-based methods including HexPlane (Cao & Johnson, 2023) and RoDynRF (Liu et al., 2023a). HexPlane represents the scene with six feature planes in both 3D space and time-space. We find that HexPlane does not produce reasonable results if only a monocular video from a single camera is given as input. Thus, other than the input monocular video, we use another video from a different viewpoint to train HexPlane for the DVS task. RoDynRF is a SoTA NeRF-based DVS method that also adopts single-view depth estimation as supervision for the 3D dynamic representations. We train it with the same single-view depth estimator GeoWizard (Fu et al., 2024) as ours. The second class is the Gaussian Splatting-based DVS methods including Deformable-GS (Yang et al., 2023) and SC-GS (Huang et al., 2024). Deformable-GS also associates a deformation field with a set of canonical Gaussians for DVS. SC-GS learns a set of keypoints and uses the deformation of these keypoints to blend the deformation of arbitrary 3D points.
@@ 4.4 Limitations
Though our method can conduct dynamic view synthesis from casually captured monocular videos, the task is still extremely challenging. One limitation is that our method can only reconstruct the visible 3D parts but cannot imagine the unseen parts, which leads to artifacts when rendering novel view videos on these unseen parts. Incorporating recent 3D-related diffusion generative models (Chung et al., 2023a; Liu et al., 2023b; Long et al., 2023) could be a promising direction to solve this problem, which we leave for future works. Another limitation is that the current training time is comparable with existing DVS methods, which could take several hours for a single scene. How to efficiently reconstruct the dynamic field would be an interesting and promising future research topic. Meanwhile, when the camera is completely static, our method strongly relies on the single-view depth estimator to estimate the 3D depth maps. Though existing single-view depth estimators (Fu et al., 2024; Ke et al., 2023; Yang et al., 2024a; Bhat et al., 2023) are trained on large-scale datasets and predict reasonable depth maps for most cases, these depth estimators may fail to capture some details which degenerate the quality. Another limitation is that we assume consistent depth order, though this assumption is satisfied by most methods (Yang et al., 2024b; Bochkovskii et al., 2024; Ke et al., 2023; Fu et al., 2024). Additionally, a challenge arises in rapidly moving scenes. Such rapid motions may result in inaccurate camera pose estimation, which is still challenging for MoDGS and all existing dynamic Gaussian reconstruction methods. Meanwhile, scenes featuring heavy specular reflections and low-light conditions cannot be well handled, which is also an extremely challenging problem in the monocular setting, which we leave for future works.
@@ 5 Conclusion
In this paper, we have presented a novel dynamic view synthesis paradigm called MoDGS. In comparison with existing DVS methods which requires “teleporting camera motions”, MoDGS is designed to render novel view images from a casually captured monocular video. MoDGS introduces two new designs to finish this challenging task. First, a new 3D-aware initialization scheme is proposed, which directly initializes the deformation field to provide a reasonable starting point for subsequent optimization. We further analyze the problem of depth loss and propose a new ordinal depth loss to supervise the learning of the scene geometry. Extensive experiments on three datasets demonstrate superior performances of our method on in-the-wild monocular videos over baseline methods.
@@ A.1 Implementation Details
We implement our MoDGS with PyTorch. To initialize the deformation field, we train it with 20k steps as stated in Sec. 3.2. Subsequently, we jointly train the 3D Gaussians and the deformation field with the rendering loss and the ordinal depth loss for another 20k steps. In Sec. 3.2, the flow is computed in evenly sampled key frames(e.g., $1/5$). And the downsampling voxel size for Gaussian initialization is $0.004^{3}$ (scenes are normalized to $[-1,1]^{3}$ ). For the outer optimization loop and rendering loss, we exactly follow the original 3DGS. And we use Gaussian centers to render depth  (Yang et al., 2023). We adopt an Adam optimizer for optimization. The learning rate for 3D Gaussians exactly follows the official implementation of 3D GS (Kerbl et al., 2023), while the learning rate of the deformation network undergoes exponential decay from 1e-3 to 1e-4 in initialization and from 1e-4 to 1e-6 in the subsequent optimization. We set $\alpha=100$ for $\ell_{\text{ordinal}}$. The weight of our depth order loss is 0.1. When computing depth ordinal loss, we first normalize the depth range to [0,1] and we only consider the depth pair with a difference larger than 0.02 for loss computation. The whole training takes around 3.5 hours to converge (2 hours for the initialization and 1.5 hours for the subsequent optimization) on an NVIDIA RTX A6000 GPU, which uses about 14G memory. The rendering speed of MoDGS is about 75 FPS.